# Promoting Constructive Deliberation: Reframing for Receptiveness

### v1

@@ Abstract
To promote constructive discussion of controversial topics online, we propose automatic reframing of disagreeing responses to signal receptiveness while preserving meaning. Drawing on research from psychology, communications, and linguistics, we identify six strategies for reframing. We automatically reframe replies according to each strategy, using a dataset of Reddit comments and replies. Through human-centered experiments, we find that the replies generated with our framework are perceived to be significantly more receptive than the original replies, as well as a generic receptiveness baseline. We analyze and discuss the implications of our results and highlight applications to content moderation. Overall, we illustrate how transforming receptiveness, a particular social science construct, into a computational framework, can make LLM generations more aligned with human perceptions.
@@ 1 Introduction
In this work, we seek to bridge NLP research on textual reframing Sharma et al. (2022); Ziems et al. (2022); Madaan et al. (2020); Chakrabarty et al. (2021) with social science theories on receptiveness. In so doing, we aim to promote pro-social deliberation online, particularly in the context of disagreement on contentious issues. We develop a computational framework to signal receptiveness while preserving the original meaning of reframed content, exploring the capabilities of LLMs for this task. We illustrate the value of translating social science theory to a computational framework by showing how model generations created with our framework better align with human notions of receptiveness.
We operationalize receptiveness through a framework of six lower-level strategies to make messages more receptive to opposing viewpoints (Section 3). We ground this conceptual work in a concrete study of Reddit posts, illustrated in Figure 1. Using an in-context framework, we compile a new corpus of automatically reframed Reddit replies. In Section 4, we assess (via automatic measures and human evaluation) the validity of the generated replies.
Section 5 then compares the reframes generated with our framework to the original replies and baselines, finding they are significantly more aligned with what humans find receptive. Reframed replies elicit relatively low negative emotional reactions, high curiosity to engage with opposing views, and reduced belief that the user is trying to shutdown discussion. We also explore the interaction between toxicity and receptiveness reframing. We find that more toxic content gets greater benefit from reframing, suggesting greater flexibility for moderation. We create a corpus of receptive-reframed replies, baseline reframes, and human annotations of receptiveness (16.6k reframes and over 9.2k total annotations) [^1] If interested, contact the authors for the dataset..
@@ 3 Approach
In Section 3.1, we introduce and motivate the six specific methods for conveying receptiveness curated from social science literature that we transform into reframing strategies. These are also illustrated in Figure 1. Following this, we describe our specific task formulation and data used (Section 3.2). Finally, we describe our method for automating this reframing (Section 3.3).
@@ 3.1 Reframing strategies for receptiveness
Brown and Levinson (1987)’s politeness theory suggests that people desire in social interactions to be appreciated and maintain a positive self image (“positive face”) and to not be curtailed or imposed upon by others (“negative face”). Disagreements may cause disrespect or offense, threatening positive face, and/or may intrude on the receiver’s autonomy, threatening negative face. Face threatening disagreements are especially evident when discussing politics on social media where partisan distrust Iyengar et al. (2019) coupled with the limited affordances available to gauge perceptions Bail (2022) may derail very well-intended interactions. Numerous strategies have been identified to mitigate perceptions of threats to face and continue the conversation while still expressing disagreement Minson and Chen (2022); Yeomans et al. (2020).
@@ Acknowledgements
includes asking for further details, clarifying understanding, and repeating what was heard. Seeking more details may be viewed as delaying expressing disagreement, potentially softening threats to negative face Pomerantz (2021). Further asking for elaboration signals expressed interest in the receiver’s view which is known to increase perceptions of receptiveness Chen et al. (2010). Examples: "what I heard you say was that…", "what did you mean by…".
Common Ground
is often an essential precondition for conflict resolution van Bijnen (2020). Highlighting and making explicit shared beliefs, assumptions and knowledge is known to increase conversational receptiveness in civic-minded workshops to reduce polarization Oliver-Blackburn and Chatham-Carpenter (2023); ListenFirst (2018). Examples: "we both believe…", "we both think…".
@@
is also called “token agreement” Pomerantz (2021) and typically takes the ‘yes,.. but’ form. This mitigation strategy allows individuals to delay and minimize the force of disagreement with a nominal token of agreement, before stating their alternate view. Examples: "I agree about …but", "I also think …yet".
@@ 3.2 Task formulation and data
Data. English Reddit data is drawn from the Debagreement dataset Pougué-Biyong et al. (2021) for disagreement detection. It consists of comments and replies from controversial subreddits: r/Brexit, r/democrats, r/Republicans, r/climate, and r/BlackLivesMatters. We exclude r/Brexit due to our US-based annotators being less familiar with Brexit-related people and issues. We select examples that are labeled as disagreement and where the comment and reply are at most 30 words each since preliminary pilots revealed that longer text increased the cognitive burden on annotators.
We run all replies through the Perspective API [^2] https://perspectiveapi.com/ and exclude any reply having predicted toxicity above 0.9, as these were typically too toxic to be reframed without significantly altering their original meaning. Further discussion of the interaction between toxicity and reframing is explored in Section 5).
@@ 3.3 Generating reframes
To automatically generate reframes for each reply, we use gpt-4 OpenAI (2023). To improve generation quality, we use in-context learning. Specifically, we provide five example comment-reply-reframe triples with each instruction prompt. The five comment-reply pairs are drawn from the Reddit data and held constant across the six reframing strategies. For each strategy and for each of the comment-reply pairs, we write a reframe of the reply for the strategy. This requires a total of 30 reframes over the six strategies and five comment-reply pairs.
Baseline reframes. We consider two zero-shot prompts with gpt-4 as baselines. For the paraphrase baseline, we prompt to simply paraphrase the reply for each comment-reply pair. This baseline serves to distinguish whether observed differences stem from our specific reframing strategies or any language model generation. For the receptiveness baseline, we directly prompt for receptiveness, without reference to any of our six strategies. This baseline allows us to evaluate our social science-informed reframing framework against a generic prompt to reframe for receptiveness. In particular, this allows us to discover whether model generations using our framework are more aligned with human perceptions of receptiveness, compared to the generic prompt. Appendix A lists full prompts and few-shot examples.
In total, we generate six reframes (plus two baseline reframes) for each of $\sim$2k comment-reply pairs in the Debagreement dataset, yielding 16.6k generations to be evaluated. We choose 450 for human validation.
@@ 4 Validation Experiments
We next present analyses conducted to validate our reframe generation methods (Section 3). We assess whether our reframes satisfy three key properties: 1) demonstrate the six distinct strategies for receptiveness; 2) preserve the meaning of original replies being reframed; and 3) remain responsive to the original comment.
@@ 4.1 Do reframes reflect the six strategies?
We inspect differences in the most common trigrams found in generated reframes compared to the original replies. We study two probability distributions: (a) $P(t|s)$, the probability of trigram $t$ being generated by strategy $s$, and (b) $P(s|t)$, given observed trigram $t$, the probability of it having been generated by strategy $s$. The former tells us whether the model is adding appropriate linguistic phrases for each strategy, while the latter measures how distinct the generations are for each strategy.
The top trigrams and probabilities are listed in Table 1. For each strategy, trigrams are shown ordered by $P(t|s)$. Regarding (a), we see that the most common trigrams being added do appear to be consistent with each strategy. In regard to (b), some strategies appear to be more unique than others. For example, Gratitude and Elaboration seem to be the most unique. While agreement, grounding, and acknowledgement share some of the most common phrases added, we note that the amount of overlap of the trigrams between the different strategies is still relatively low (Table 5 in Appendix B).
@@ 4.2 Do the reframes preserve meaning?
Table 2: Automatic validation of meaning preservation of model generations. Distinct ngrams and SBERT score are taken between original dialogue (comment + reply) and reframed reply. PC indicates the proportion of contradictions between original and reframed replies. Values for all columns are on a [0,1] scale. $\downarrow$ / $\uparrow$ arrows indicate that smaller/larger values are better. “B” indicates a baseline.
@@ 4.3 Are the reframes contextually relevant?
We evaluate whether our reframing framework produces replies that actually reply to the comment, (i.e., that are contextually relevant). For this analysis, we compare our reframes with the original reply from Reddit. We ask annotators whether each reply shown constitutes a reasonable response to the original comment using a 5-point Likert scale (0 = reply doesn’t make sense, 4 = the reply is very reasonable) in order to determine whether annotators think a reply is relevant in context of the comment. The interface is shown in Appendix Figure 4.
Annotations are collected via Amazon Mechanical Turk (MTurk). To ensure quality annotation, we apply two strategies to curate our annotator pool: (1) sourcing workers already pre-qualified by Amazon for its Ground Truth workforce [^4] https://aws.amazon.com/sagemaker/groundtruth/, and (2) filtering other workers ourselves, requiring at least 97% accuracy and 5000 completed hits. We also require workers to pass a qualification test consisting of four examples similar to our main task. Finally, given the US-centric nature of our data, we require workers to be US-based. We pay workers \$10/hr. We give annotators 420 examples in a multiple batches, and we assign 3-7 workers per example (3.7 on average). Reasonability scores are averaged across annotators and examples within a strategy. Table 3 shows results.
@@
We conduct a paired t-test between the examples for each strategy and the original to determine if the difference is statistically significant. We find that the mean reasonability score for each of the strategies is significantly higher than the reasonability of the original reply (see Table 10 in Appendix E for mean differences, test statistics, and p-values). This indicates that our reframes are perceived to be reasonable replies to the original comments.
Surprisingly, the original, human-written replies are perceived to be significantly less reasonable, likely because some of the original replies do not directly respond to the comment (as is common for social media dialogue), while the generated replies almost always take the context of the original comment into account. See Table 8 in Appendix C for examples. Another possible reason is that annotators may have interpreted reasonable response to mean more receptive which may have affected their ratings.
@@ 5.1 Annotation task
We conduct an experiment on MTurk to evaluate whether the generated reframes are perceived to be more receptive than the original replies. Section 4.3 described earlier how we filter the MTurk workforce to ensure quality of collected data.
As discussed in Section 3.2, our data consists of comment-reply pairs, along with generated reframes for two baselines and our receptiveness framework. We use 9 versions of each of the 75 original replies: the original, the two baselines, and the six strategy reframes. Annotators are asked to compare the relative receptiveness between two alternative replies to a comment. One reply is the original and the other is either one of our reframes or a baseline. Annotation interfaces are shown in Figures 5, 6, and 7 in Appendix F. Annotation is performed in multiple batches with 3-6 workers per example (4.4 on average). To evaluate the reframes, we develop a receptiveness index translated from social science literature, as described below.
@@ 5.2 Dependent variable: Receptiveness index
We translate Minson et al. (2020)’s receptiveness index for dispositional receptiveness to conversational receptiveness. Their index was determined via an 18 dispositional statement scale. Based on participant responses, they surfaced a set of four factors that underlie these 18 statements:
@@
To translate the scale to the fast-paced MTurk platform and to our conversational receptiveness task, we reduce the quantity and transform two statements into two questions per factor (Table 9 in Appendix D). The questions compare the receptiveness level of the reframed or baseline reply vs. the original reply, though annotators are not told which is which. We use a 7-point Likert scale for each question: -3 indicates the original reply is much more receptive, +3 indicates the opposite, and 0 indicates that original and reframed reply are equally receptive. Answers are averaged over the eight questions (and over all annotations) to determine our receptiveness index for each example.
@@ 5.3 Results
To compare mean receptiveness scores between our reframed replies vs. baseline replies, we use a mixed effects logistic regression model via the lme4 package Bates et al. (2015). We model the receptiveness score with a random effect for the comment ID and fixed effect for the strategy used.
We show the regression co-efficients in Appendix Table 11. The estimated marginal mean and the 95% CI receptiveness scores of the strategies and the baseline are shown in Figure 2. A marginal mean of zero implies the reframed reply is no different from the original reply with respect to receptiveness. From Figure 2, we find that neither the strategies nor baselines overlap with 0, indicating that they are more receptive than the original. The reframes using our strategies have a statistically significant higher receptiveness index than both baselines. This indicates that annotators consider generations from our strategies more receptive than the paraphrases or baseline receptive reframes. In other words, the replies generated using our social science-informed in-context framework are more aligned with human notions of receptiveness, compared to a generic prompting approach. The mean differences, test statistics and p-values are in Table 12 in Appendix E.
@@ 5.4 Factor breakdown of receptiveness
We have seen that the framework-generated reframes are more aligned with human perceptions of receptiveness, but why is this the case? We investigate this by examining the receptiveness scores split by the individual factors. To do so, we averaged the values for the two questions in each factor. We can see from Table 4 that while our reframes appear to decrease negative emotions, increase openness, and increase curiosity towards opposing views, they interestingly do not do as well in reducing bias towards the opposing party. We discuss this further in Section 6. We note that such a factor analysis (drawing upon social science theory) provides deeper insight into the reasoning for the receptiveness scores than would be possible with the direct receptive baseline.

### v2

@@ Abstract
To promote constructive discussion of controversial topics online, we propose automatic reframing of disagreeing responses to signal receptiveness to a preceding comment. Drawing on research from psychology, communications, and linguistics, we identify six strategies for reframing. We automatically reframe replies to comments according to each strategy, using a Reddit dataset. Through human-centered experiments, we find that the replies generated with our framework are perceived to be significantly more receptive than the original replies and a generic receptiveness baseline. We illustrate how transforming receptiveness, a particular social science construct, into a computational framework, can make LLM generations more aligned with human perceptions. We analyze and discuss the implications of our results, and highlight how a tool based on our framework might be used for more teachable and creative content moderation.
@@ 1 Introduction
In this work, we seek to bridge NLP research on textual reframing Sharma et al. (2022); Ziems et al. (2022); Madaan et al. (2020); Chakrabarty et al. (2021) with social science theories on receptiveness. Our aim is to promote pro-social deliberation online; particularly to foster productive discussion for contentious issues. We develop a computational framework to signal receptiveness while preserving the original meaning of reframed content, studying the capabilities of LLMs for this task. We illustrate the value of translating social science theory to a computational framework by showing how model generations created with our framework better align with human notions of receptiveness, allowing for more useful content moderation tools.
Models that are more receptive-aligned might allow for content moderation tools that are valuable to users and moderators despite being automated. For moderators, a tool built on our receptiveness framework could allow for more scalable and creative moderation. For users, such a tool might provide learning opportunities, suggesting possible revisions if their post is flagged as unreceptive. Users could learn how to be more receptive from suggestions whether or not they accept them.
Here, we operationalize receptiveness through a framework of six lower-level strategies from social science theory to make messages more receptive to opposing viewpoints. We ground this conceptual work in a concrete study of Reddit posts, illustrated in Figure 1. We compile a new corpus of automatically reframed Reddit replies and assess (via automatic measures and human evaluation) the validity of the generated replies.
We compare reframes generated by our framework to original replies and baselines. We find they are significantly more aligned with what humans find receptive, eliciting relatively low negative emotional reactions, high curiosity to engage with opposing views, and reduced belief that the user is trying to shut down discussion. We also explore the interaction between toxicity and receptiveness reframing, finding that more toxic content benefits more from reframing. We create and share a corpus of receptive-reframed replies, baseline reframes, and human annotations of receptiveness (16.6k reframes and over 9.2k total annotations) [^1] Contact authors for dataset if interested.
@@ 3 Approach
In Section 3.1, we introduce and motivate six specific methods we apply for conveying receptiveness curated from social science literature that we transform into reframing strategies. These are also illustrated in Figure 1. Following this, we describe our specific task formulation and data used (Section 3.2). Finally, we describe our method for automating this reframing (Section 3.3).
@@ 3.1 Reframing strategies for receptiveness
Brown and Levinson (1987)’s politeness theory suggests that people desire in social interactions to be appreciated and maintain a positive self image (“positive face”) and to not be curtailed or imposed upon by others (“negative face”). Disagreements may cause disrespect or offense, threatening positive face, and/or may intrude on the receiver’s autonomy, threatening negative face. Face threatening disagreements are especially evident when discussing politics on social media where partisan distrust Iyengar et al. (2019) coupled with the limited affordances available to gauge perceptions Bail (2022) may derail even well-intended interactions. Numerous strategies have been identified to mitigate perceptions of threats to face and continue the conversation while still expressing disagreement Minson and Chen (2022); Yeomans et al. (2020).
@@ Acknowledgements
includes asking for further details, clarifying understanding, and repeating what was heard. Seeking more details may be viewed as delaying expressing disagreement, potentially softening threats to negative face Pomerantz (2021). Further asking for elaboration signals expressed interest in the speaker’s view which is known to increase perceptions of receptiveness Chen et al. (2010). Examples: "what I heard you say was that…", "what did you mean by…".
Grounding
is often an essential precondition for conflict resolution van Bijnen (2020). Highlighting and making explicit shared beliefs, assumptions and knowledge is known to increase conversational receptiveness in civic-minded workshops to reduce polarization Oliver-Blackburn and Chatham-Carpenter (2023); ListenFirst (2018). While acknowledgement simply conveys understanding of the speaker’s point of view, grounding emphasizes common ground and commonalities between the discussants’ viewpoints. Examples: "we both believe…", "we both think…".
@@
typically takes the ‘yes,.. but’ form Pomerantz (2021). This mitigation strategy allows individuals to delay and minimize the force of disagreement with an explicit but nominal token of agreement. This strategy can be viewed as a narrow subset of the grounding strategy. While agreement simply makes explicit some agreement about the speaker’s statements, grounding also includes highlighting even broader shared beliefs (for example, “I am also a libertarian but I disagree on this …”). We separate out this strategy from grounding as our expectation is that highlighting specific agreement with the speaker’s statements may signal more receptiveness than emphasizing broader common ground. Examples: "I agree about …but", "I also think …yet".
@@ 3.2 Task formulation and data
Data. English Reddit data is drawn from the Debagreement dataset Pougué-Biyong et al. (2021) for disagreement detection. It consists of comments and replies from controversial subreddits: r/Brexit, r/democrats, r/Republicans, r/climate, and r/BlackLivesMatters. We exclude r/Brexit due to our US-based annotators being less familiar with Brexit-related people and issues. We select examples that are labeled as disagreement and where the comment and reply are at most 30 words each, since preliminary pilots revealed that longer text increased the cognitive burden on annotators. The average word length of the comments and replies in the Reddit data are 36.15 words and 33.12 words respectively, indicating our length-constrained data is not much smaller than average.
We run all replies through the Perspective API [^2] https://perspectiveapi.com/ and exclude any reply with predicted toxicity above 0.9, as these were typically too toxic to be reframed without significantly altering their original meaning. Further discussion of the interaction between toxicity and reframing is explored in Section 5).
@@ 3.3 Generating reframes
To generate reframes for each reply, we use gpt-4 OpenAI (2023). To improve generation quality, we use in-context learning. Specifically, we provide five example comment-reply-reframe triples with each instruction prompt. The five comment-reply pairs are drawn from the Reddit data and held constant across the six reframing strategies. For each strategy and for each of the comment-reply pairs, we write a reframe of the reply for the strategy. This requires a total of 30 reframes over the six strategies and five comment-reply pairs.
Baseline reframes. We consider two zero-shot prompts with gpt-4 as baselines. For the paraphrase baseline, we prompt to simply paraphrase the reply for each comment-reply pair. This baseline serves to distinguish whether observed differences stem from our specific reframing strategies or any language model generation. For the receptiveness baseline, we directly prompt for receptiveness, without reference to any of our six strategies. This baseline allows us to evaluate our social science-informed reframing framework against a generic prompt to reframe for receptiveness.
In total, we generate six reframes (plus two baseline reframes) for each of $\sim$2k comment-reply pairs in the Debagreement dataset, yielding 16.6k generations. Appendix B lists full prompts and few-shot examples.
@@ 4 Validation Experiments
We next present analyses conducted to validate our reframe generation methods. We assess whether our reframes satisfy three key properties: 1) demonstrate the six distinct strategies for receptiveness; 2) preserve the meaning of original replies being reframed; and 3) remain responsive to the original comment.
@@ 4.1 Do reframes reflect the six strategies?
We inspect differences in the most common trigrams found in generated reframes vs. the original replies. We study two probability distributions: (a) $P(t|s)$, the probability of trigram $t$ being generated by strategy $s$, and (b) $P(s|t)$, given observed trigram $t$, the probability of it having been generated by strategy $s$. The former tells us whether the model is adding appropriate linguistic phrases for each strategy, while the latter measures how distinct the generations are for each strategy.
The top trigrams and probabilities are listed in Table 1. For each strategy, trigrams are shown ordered by $P(t|s)$. Regarding (a), we see that the most common trigrams being added do appear to be consistent with each strategy. Further, except for the “thank you for” trigram, trigrams appear less than 10% of the time as part of each strategy, implying that there is reasonable variety in the way the strategy is expressed. In regard to (b), we find that the majority of the strategies are unique. However, the top trigrams in agreement overlap with grounding and acknowledgement. Recall that agreement represents a narrower subset of grounding; while agreement simply makes explicit any agreement with the speaker’s statements, grounding also includes highlighting broad shared beliefs. Further, acknowledgement conveys understanding of the speaker’s statement without necessarily agreeing with them. While the overlap of the “i agree that” trigram with grounding is expected, overlaps with acknowledgement are rather surprising and suggest that better prompts for agreement were needed. We discuss more about this issue in the limitations section. The amount of trigram overlap between the different strategies is shown in Appendix Figure 4.
@@ 4.2 Do the reframes preserve meaning?
Table 2: Automatic validation of meaning preservation of model generations. Distinct ngrams and SBERT score are taken between original dialogue (comment + reply) and reframed reply. PC indicates the proportion of contradictions between original and reframed replies. All values lie in [0,1]. $\downarrow$ / $\uparrow$ arrows mean smaller/larger values are better. “B” indicates a baseline.
@@ 4.3 Are the reframes contextually relevant?
We evaluate whether our reframing framework produces replies that are contextually relevant to the comment. For this analysis, we compare a subset of our reframes with the original reply from Reddit. We ask annotators whether each reply shown constitutes a reasonable response to the original comment using a 5-point Likert scale (0 = reply doesn’t make sense, 4 = the reply is very reasonable) to determine whether annotators think a reply is relevant in context of the comment. The interface is shown in Appendix Figure 5. Annotations are collected via Amazon Mechanical Turk (MTurk). We describe the steps taken to ensure quality annotation in Appendix A. We give annotators 420 examples in multiple batches, and we assign 3-7 workers per example (3.7 on average). Reasonability scores are averaged across annotators and examples within a strategy. Table 3 shows results.
@@
We conduct a paired t-test between the examples for each strategy and the original to determine if the difference is statistically significant. We find that the mean reasonability score for each of the strategies is significantly higher than the reasonability of the original reply (see Appendix Table 9 for mean differences, test statistics, and p-values). This indicates that our reframes are perceived to be reasonable replies to the original comments.
Surprisingly, the original, human-written replies are perceived to be significantly less reasonable, likely because some of the original replies do not directly respond to the comment (common in social media), while the generated replies almost always take the context of the original comment into account. See Appendix Table 7 for examples. Another possible reason is that annotators may have interpreted reasonable response to mean more receptive which may have affected their ratings.
@@ 5.1 Annotation task
We conduct an experiment on MTurk to evaluate whether the generated reframes are perceived to be more receptive than the original replies. Appendix A describes how we filter the MTurk workforce to ensure quality of collected data.
As discussed in Section 3.2, our data consists of comment-reply pairs, along with generated reframes for two baselines and our receptiveness framework. We use 9 versions of each of the 75 original replies: the original, the two baselines, and the six strategy reframes. Annotators are asked to compare the relative receptiveness between two alternative replies to a comment: the original vs. one of our reframes or a baseline generation. Annotation interfaces are shown in Figures 6, 7, and 8 in Appendix G. Annotation is performed in multiple batches with 3-6 workers per example (4.4 on average). Given the subjective nature of this annotation task, inter-annotator agreement is rather low, yielding a Krippendorff’s $\alpha$ score of 0.23. Given this low agreement, we study the 95% confidence interval around the receptiveness scores to evaluate how this disagreement affects variance around the mean receptiveness scores. To evaluate the reframes, we develop a receptiveness index translated from social science literature, as described below.
@@ 5.2 Dependent variable: Receptiveness index
We translate Minson et al. (2020)’s receptiveness index for dispositional receptiveness to conversational receptiveness. Their index was quantified via an 18 dispositional statement scale. Based on responses, they surfaced a set of four factors:
@@
To translate the scale to the fast-paced MTurk platform and to our conversational receptiveness task, we reduce the quantity and transform the statements into two questions per factor (Appendix Table 8). The questions compare the receptiveness of the reframed or baseline reply to the original reply, though annotators are not told which is which. We use a 7-point Likert scale for each question: -3 indicates the original reply is much more receptive, +3 indicates the opposite. Answers are averaged over the eight questions to determine our receptiveness index for each example.
@@ 5.3 Results
To compare mean receptiveness scores between our reframed and baseline replies, we use a mixed effects logistic regression model via the lme4 package Bates et al. (2015). We model the receptiveness score with a random effect for the comment ID and fixed effect for the strategy used.
We show the regression co-efficients in Appendix Table 10. The estimated marginal mean and the 95% CI receptiveness scores of the strategies and the baseline are shown in Figure 2. A marginal mean of zero implies the reframed reply is no different from the original reply with respect to receptiveness. From Figure 2, we find that neither the strategies nor baselines overlap with 0, indicating that they are more receptive than the original. The reframes using our strategies have a statistically significant higher receptiveness index than both baselines. This indicates that annotators consider generations from our strategies more receptive than the paraphrases or baseline receptive reframes. In other words, the replies generated using our social science-informed framework are more aligned with human notions of receptiveness, compared to a generic prompting approach. The mean differences, test statistics and p-values are in Appendix Table 11.
@@ 5.4 Factor breakdown of receptiveness
We have seen that the framework-generated reframes are more aligned with human perceptions of receptiveness, but why is this the case? We investigate this by examining the receptiveness scores split by the individual factors. To do so, we averaged the values for the two questions in each factor. We can see from Table 4 that while our reframes appear to decrease negative emotions, increase openness, and increase curiosity towards opposing views, they interestingly do not do as well in reducing bias towards the opposing party. We discuss this further in Section 6. Such a factor analysis (drawing upon social science theory) provides deeper insight into the reasoning for the receptiveness scores than would be possible with the direct receptive baseline.